OpenStudy (anonymous):

integrate t^2dt/(1+t^2)^1/2

ganeshie8 (ganeshie8):

familiar with hyperbolic functions ?

ganeshie8 (ganeshie8):

\[\large \int \dfrac{dx}{\sqrt{1+x^2}} = \sin h ^{-1}(x)+C\]

ganeshie8 (ganeshie8):

seen that before ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

Okay, then drop that idea, and you may work it using trig substitution

ganeshie8 (ganeshie8):

look at the stuff inside radical in the bottom

ganeshie8 (ganeshie8):

which trig substitution is appropriate here ?

OpenStudy (anonymous):

tan

ganeshie8 (ganeshie8):

do it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am not getting

ganeshie8 (ganeshie8):

you may use tables also : \[ \large \begin{align} \int \dfrac{t^2}{\sqrt{1+t^2}}~dt &= \int \dfrac{(1+t^2)-1}{\sqrt{1+t^2}}~dt \\~\\&= \int \sqrt{1+t^2}~dt ~-~\int \dfrac{1}{\sqrt{1+t^2}}~dt \end{align}\]

ganeshie8 (ganeshie8):

look up in your formulas, you may find some formula for \(\large \int \sqrt{1+x^2} dx\) otherwise use trig substitution and derive the formulas now

OpenStudy (anonymous):

i got it thanks a lot

ganeshie8 (ganeshie8):

#29 and #32 formulas http://integral-table.com/

ganeshie8 (ganeshie8):

yw :)